an executive invest $22000, some at 7% and some at6% annual interest. If he recieves an annual return of $1420, how much is invested in each rate?: an executive invest $22000, some at 7% and some at6% annual interest. If he recieves an annual return of $1420, how much is invested in each rate?
Answer by checkley77(3394)

.07X+.06(22,000-X)=1,420
.07X+1,320-.06X=1,420
.02X=1,420-1,320
.01X=100
X=100/.01
X=$10,000 INVESTED @ 7%.
22,000-10,000=$12,000 INVESTED @ 6%.
PROOF
.07*10,000+.06*12,000=1,420
700+720=1,420
1,420=1,420